You are considering investing RM68000 in new equipment. You estimate that the net cash flows wll be RM12000 during the first year, but will increase by RM2500 per year the next year and each year thereafter. The equipment is estimated to have a 1-year service life and a net salvage value ofRM5200 at that time. Assume MARR of 6%. a Calculate the annual capital cost CR (ownership cost) for the equipment. b.Determine the equivalent annual savings. c.Is this a wise investment? YN Format: 7392 Format: 49027 Format A Submit Pev 1234
Answer #1

a) Annual capital cost or Annual worth (Costs) for the equipment =68000+12000(A/F, 6%,1) + 2500(A/G,6%,10)

                                                                           =68000 +12000* 1.000 +2500* 4.0220 = RM90055

b) Net salvage value at the end of service life of equipment = RM5200

Equivalent Annual savings or Annual worth (benefits) for the equipment = 5200(F/A,6%,11) = 5200* 14.9716 = RM77852.32

c) Benefit cost or B/C ratio of the equipment = AW(Benefits)/AW(costs) = 77852.32/90055= 0.86<1

Therefore, the investment is not wise- N